Let f(x)= x^2-25 and g(x)=x+5. Find and simplify (f/g)(x). State the domain of (f/g)(x).
Could you please show me all the work that is necessary to complete this problem so that I understand it completely and will know how it was done.

Answer by rothauserc(4718)

(f/g)(x) = (x^2 -25) / (x +5)
now factor the numerator
(f/g)(x) = ((x +5) * (x -5)) / (x +5)
now we can cancel (x +5) in the numerator and denominator
(f/g)(x) = x -5
domain of (f/g)(x) is the real numbers